# Learning in Wilson-Cowan model for metapopulation

### v1

@@ 5.2.1 Wilson-Cowan model for metapopulation in combination with a Convolutional Neural Network
We then decided to implement a VGG-16 architecture (see A.2) for classifying the CIFAR-10 dataset. In this case, we observed that the accuracy obtained is on average $\psi=86.59\%$ (with the best performance at $\psi_{best}=86.83\%$), thanks to the deep convolutional architecture used. Although our model lacks specific dropout and batch normalization layers, and data augmentation techniques that are typically used in the best VGG-16 implementations, our performance remains reasonably high compared to many other CNN architectures. For further comparison, the reader shall refer to the performance benchmarks available at this link (Meta-AI-Research, 2024).
For the last analysis, we chose to check if transfer learning still works with our model. In general, transfer learning can be used to improve performance on a task A for which training data is in short supply by using data from a related task B, for which data is more plentiful. The two tasks should have the same kind of inputs, and there should be some commonality between the tasks so that low-level features, or internal representations, learned from task B will be useful for task A. When data for task A is very scarce, we might simply retrain the final layer of the network. In contrast, if there are more data points, it is feasible to retrain several layers. This process of learning parameters using one task that are then applied to one or more other tasks is called pre-training (Bishop and Bishop, 2023).
Figure 10: Description of TF-FLOWERS dataset. The figure illustrates the composition of TF-FLOWERS dataset. The first row shows the labels of the examples found in the second and the third row. The fourth row presents the five target images we created to serve as attractors for the dynamics.
We, therefore, confined ourselves to work with a small dataset, TF-FLOWERS (see reftfflowers for details). This dataset is composed of $3670$ colored $224\times 224\times 3$ images of flowers, categorized into $K=5$ distinct classes: daisy, dandelion, roses, sunflowers, and tulips. An example of such images is presented in Fig. 10. It is our task A, and as the reader can see, the dataset is scarce in the number of images, but not in the details of each image. We, therefore, decided to use the VGG16 model trained on the IMAGENET dataset [^4] The IMAGENET dataset (Deng et al., 2009) comprises $14$ million natural images each of which has been hand labelled into one of nearly $22000$ categories. A subset of images comprising $1000$ non-overlapping categories is often used for pre-training models in deep learning. The fact to have so many categories made the problem much more challenging because, if the classes were distributed uniformly, random guessing would have an error rate of $99.9\%$. (task B) for transfer learning, and then train just the last layers, as well as our Wilson-Cowan model for metapopulation (see A.4).
@@ 5.2.2 Wilson-Cowan model for metapopulation in combination with a Pre-Trained Transformer
We have previously observed that the use of CNNs significantly enhances the accuracy of models. These networks draw inspiration from biological systems and exhibit remarkable similarities to specific regions of the human brain. Recently, CNNs have also been employed to describe and understand how emotional patterns are integrated into the human visual system (Kragel et al., 2019). However, CNNs fall short in replicating another fundamental capability of the human brain: the comprehension and production of language.
Human language processing is one of the most fascinating capabilities of our brain. It relies on a set of interconnected brain areas in the frontal and temporal lobes, typically in the left hemisphere, forming a network. This language network supports both comprehension (spoken, written, and signed) (Deniz et al., 2019; Terri L. Scott and Fedorenko, 2017) and production (Menenti et al., 2011; Hu et al., 2022). It has been extensively studied to understand its sensitivity to linguistic regularities at multiple levels (Regev et al., 2024; Fedorenko et al., 2020).
@@
Transformers, as described by Vaswani et al. (2017b) in their seminal work, have become the most trending topic in natural language processing (NLP) due to their outstanding performance in capturing formal linguistic competence—i.e., the knowledge of rules and statistical regularities of language. However, they exhibit limitations in functional linguistic competence, which involves the practical use of language in real-world situations (Mahowald et al., 2024).
This section explores the potential synergy between our model and a transformer architecture. We explore this synergy to address the limitations of the Wilson-Cowan model for metapopulation in extrapolating the complex structural patterns embedded in language. Therefore, we utilize a transformer architecture, specifically BERT (Devlin et al., 2018), to extract the most important features from the text for a classification task with our Neural Mass Network Model.
@@ 6 Conclusion
However, backpropagation can be viewed as an efficient way to achieve reasonable parameter estimates, which can then be subjected to further testing. Even if backpropagation is considered merely a technical solution, the trained model may still serve as a good approximation of neural systems. Currently, many researchers are exploring new supervised learning optimization algorithms that are more biologically valid, i.e., neurobiologically plausible methods by which the brain could adjust its internal parameters to optimize objective functions (Song et al., 2024). Future publications will aim to validate new biologically plausible Wilson-Cowan models for metapopulations that can effectively learn visual or textual patterns.
@@ A.1 MNIST and Fashion MNIST
The MNIST dataset (Deng, 2012) is composed by $70000$ grey scale handwritten images of size $28\times 28$ in $K=10$ classes. The dataset is divided in $60000$ images for training set and $10000$ for test set.
The Fashion-MNIST dataset (Xiao et al., 2017) is composed by $70000$ grey scale Zalando’s article images of size $28\times 28$ in $K=10$ classes. The dataset is divided in $60000$ images for training set and $10000$ for test set.
@@ A.2 VGG-16
The VGG-16 (Simonyan and Zisserman, 2015) model, where VGG stands for the Visual Geometry Group, who developed the model, and 16 refers to the number of learnable layers in the model, has some simple designed principles leading to a relative uniform architecture that minimizes the number of hyperparamete choices that need to be made. In principle, it was developed to take an input image having $224\times 224\times 3$ colored pixels (RGB channels), followed by sets of convolutional and pooling layers for downsampling. It can be also applied to smaller color images $32\times 32\times 3$. Here we present the original architecture as described in (Bishop and Bishop, 2023).
@@ A.3 CIFAR10
The CIFAR-10 dataset (Krizhevsky et al., 2009) is composed by $60000$ colored images of size $32\times 32\times 3$ in $K=10$ classes (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck), with $6000$ images per class. There are $50000$ training images and $10000$ test images. The CNN for this analysis is composed by the VGG16 architecture (see A.2). After these layers, a flatten layer is used to convert the output for the dense layers. The CNN output is then passed to three distinct dense layers with $2048$, $1024$, and $784$ neurons, all using the ReLU activation function. The second and third dense layers are preceded by a batch normalization layer. The output of the third dense layer is then passed to our Wilson-Cowan model for metapopulation, with $\mathcal{N}=784$. The total number of parameters is $19294817$, divided into $615441$ parameters for our Wilson-Cowan model for metapopulation and $18679376$ for the CNN part plus the three dense layers.
@@ A.4 TF-FLOWERS
The TF-FLOWERS dataset (TensorFlow-Team, 2019) is composed by $3670$ colored $224\times 224\times 3$ images of flowers, categorized into $K=5$ distinct classes: daisy, dandelion, roses, sunflowers, and tulips. We chose $90\%$ of the dataset as training images and $10\%$ test images. The CNN for this analysis is composed by the VGG16 architecture (see A.2), with pre-trained weights given by IMAGENET dataset. After these layers, a flatten layer is used to convert the output for the dense layers. The CNN output is then passed to four distinct dense layers with $4096$, $2048$, $1024$, and $784$ neurons, all using the ReLU activation function. The second and third dense layers are preceded by a batch normalization layer. The output of the third dense layer is then passed to our Wilson-Cowan model for metapopulation, with $\mathcal{N}=784$. The total number of parameters is $129460641$, divided into $615441$ parameters for our Wilson-Cowan model for metapopulation and $128845200$ for the CNN part plus the three dense layers.
@@ A.5 IMDB
The IMDB (Internet Movie Database) dataset (Pal et al., 2020) is a comprehensive and widely-used dataset in the field of machine learning and data analysis, particularly for tasks involving natural language processing (NLP) and sentiment analysis. This dataset contains extensive information on movies, television shows, and other forms of visual entertainment. The labeled data set consists of $50000$ IMDB movie reviews, specially selected for sentiment analysis. The sentiment of reviews is binary, meaning the IMDB rating $<5$ results in a sentiment score of 00, and rating $\geq 7$ have a sentiment score of $1$. No individual movie has more than $30$ reviews. The $25000$ review labeled training set does not include any of the same movies as the $25000$ review test set. In addition, there are another $50000$ IMDB reviews provided without any rating labels.
@@
This language model is then associated with a dropout layer with a parameter $p=0.5$ and a dense layer of $512$ neurons with a sigmoid activation function. Following this, we integrate our Wilson-Cowan model for metapopulation (with $\mathcal{N}=512$), ending up with a final model with $110138626$ As stated in the main text, we perform a modification for this particular classification task. Specifically, we do not compare the final state of the dynamics with the respective planted eigenvector anymore. Instead, we apply a non-linear transformation to the final state, normalizing the inverse of equation (11), to obtain a straightforward output for our $K=2$ class problem. Consequently, we derive a probability to be in one of the two classes, allowing us to apply binary cross-entropy as the loss function and utilize binary accuracy to test the performance of our learning algorithm. We used a learning rate of $0.00003$, and we fixed the number of epochs for the fine-tuning of the whole network at $10$, with mini-batch size equal to $4$. The final time $T$ was set to be equal $4.0\Delta t^{-1}$, with $\Delta t=0.1$. For the tokenization of the dataset, we have followed the tutorial given at this link(TensorFlow, 2024).

### v2

@@ 5.2.1 Wilson-Cowan model for metapopulation in combination with a Convolutional Neural Network
We then decided to implement a VGG-16 architecture (see A.2) for classifying the CIFAR-10 dataset. In this case, we observed that the accuracy obtained is on average $\psi=86.59\%$ (with the best performance at $\psi_{best}=86.83\%$), thanks to the deep convolutional architecture used. Although our model lacks specific dropout and batch normalization layers, and data augmentation techniques that are typically used in the best VGG-16 implementations, our performance remains reasonably high compared to many other CNN architectures. For further comparison, the reader shall refer to the performance benchmarks available at this link 444 https://paperswithcode.com/sota/image-classification-on-cifar-10 .
For the last analysis, we chose to check if transfer learning still works with our model. In general, transfer learning can be used to improve performance on a task A for which training data is in short supply by using data from a related task B, for which data is more plentiful. The two tasks should have the same kind of inputs, and there should be some commonality between the tasks so that low-level features, or internal representations, learned from task B will be useful for task A. When data for task A is very scarce, we might simply retrain the final layer of the network. In contrast, if there are more data points, it is feasible to retrain several layers. This process of learning parameters using one task that are then applied to one or more other tasks is called pre-training [10].
Figure 10: Description of TF-FLOWERS dataset. The figure illustrates the composition of TF-FLOWERS dataset. The first row shows the labels of the examples found in the second and the third row. The fourth row presents the five target images we created to serve as attractors for the dynamics ($\mathcal{N}=784$).
We, therefore, confined ourselves to work with a small dataset, TF-FLOWERS (see reftfflowers for details). This dataset is composed of $3670$ colored $224\times 224\times 3$ images of flowers, categorized into $K=5$ distinct classes: daisy, dandelion, roses, sunflowers, and tulips. An example of such images is presented in Fig. 10. It is our task A, and as the reader can see, the dataset is scarce in the number of images, but not in the details of each image. We, therefore, decided to use the VGG16 model trained on the IMAGENET dataset [^5] The IMAGENET dataset [22] comprises $14$ million natural images each of which has been hand labelled into one of nearly $22000$ categories. A subset of images comprising $1000$ non-overlapping categories is often used for pre-training models in deep learning. The fact to have so many categories made the problem much more challenging because, if the classes were distributed uniformly, random guessing would have an error rate of $99.9\%$. (task B) for transfer learning, and then train just the last layers, as well as our Wilson-Cowan model for metapopulation (see A.4).
@@ 5.2.2 Wilson-Cowan model for metapopulation in combination with a Pre-Trained Transformer
We have previously observed that the use of CNNs significantly enhances the accuracy of models. These networks draw inspiration from biological systems and exhibit remarkable similarities to specific regions of the human brain. Recently, CNNs have also been employed to describe and understand how emotional patterns are integrated into the human visual system [49]. However, CNNs fall short in replicating another fundamental capability of the human brain: the comprehension and production of language.
Human language processing is one of the most fascinating capabilities of our brain. It relies on a set of interconnected brain areas in the frontal and temporal lobes, typically in the left hemisphere, forming a network. This language network supports both comprehension (spoken, written, and signed) [24] and production [39]. It has been extensively studied to understand its sensitivity to linguistic regularities at multiple levels [80].
@@
Transformers, as described by [92] in their seminal work, have become the most trending topic in natural language processing (NLP) due to their outstanding performance in capturing formal linguistic competence—i.e., the knowledge of rules and statistical regularities of language. However, they exhibit limitations in functional linguistic competence, which involves the practical use of language in real-world situations [56].
This section explores the potential synergy between our model and a transformer architecture. We explore this synergy to address the limitations of the Wilson-Cowan model for metapopulation in extrapolating the complex structural patterns embedded in language. Therefore, we utilize a transformer architecture, specifically BERT [25], to extract the most important features from the text for a classification task with our Neural Mass Network Model.
@@ 6 Conclusion
However, backpropagation can be viewed as an efficient way to achieve reasonable parameter estimates, which can then be subjected to further testing. Even if backpropagation is considered merely a technical solution, the trained model may still serve as a good approximation of neural systems. Currently, many researchers are exploring new supervised learning optimization algorithms that are more biologically valid, i.e., neurobiologically plausible methods by which the brain could adjust its internal parameters to optimize objective functions [87]. Future publications will aim to validate new biologically plausible Wilson-Cowan models for metapopulation that can effectively learn visual or textual patterns.
@@ A.1 MNIST and Fashion MNIST
The MNIST dataset [23] is composed by $70000$ grey scale handwritten images of size $28\times 28$ in $K=10$ classes. The dataset is divided in $60000$ images for training set and $10000$ for test set.
The Fashion-MNIST dataset [97] is composed by $70000$ grey scale Zalando’s article images of size $28\times 28$ in $K=10$ classes. The dataset is divided in $60000$ images for training set and $10000$ for test set.
@@ A.2 VGG-16
The VGG-16 [86] model, where VGG stands for the Visual Geometry Group, who developed the model, and 16 refers to the number of learnable layers in the model, has some simple designed principles leading to a relative uniform architecture that minimizes the number of hyperparamete choices that need to be made. In principle, it was developed to take an input image having $224\times 224\times 3$ colored pixels (RGB channels), followed by sets of convolutional and pooling layers for downsampling. It can be also applied to smaller color images $32\times 32\times 3$. Here we present the original architecture as described in [10].
@@ A.3 CIFAR10
The CIFAR-10 dataset [51] is composed by $60000$ colored images of size $32\times 32\times 3$ in $K=10$ classes (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck), with $6000$ images per class. There are $50000$ training images and $10000$ test images. The CNN for this analysis is composed by the VGG16 architecture (see A.2). After these layers, a flatten layer is used to convert the output for the dense layers. The CNN output is then passed to three distinct dense layers with $2048$, $1024$, and $784$ neurons, all using the ReLU activation function. The second and third dense layers are preceded by a batch normalization layer. The output of the third dense layer is then passed to our Wilson-Cowan model for metapopulation, with $\mathcal{N}=784$. The total number of parameters is $19294817$, divided into $615441$ parameters for our Wilson-Cowan model for metapopulation and $18679376$ for the CNN part plus the three dense layers.
@@ A.4 TF-FLOWERS
The TF-FLOWERS dataset [91] is composed by $3670$ colored $224\times 224\times 3$ images of flowers, categorized into $K=5$ distinct classes: daisy, dandelion, roses, sunflowers, and tulips. We chose $90\%$ of the dataset as training images and $10\%$ test images. The CNN for this analysis is composed by the VGG16 architecture (see A.2), with pre-trained weights given by IMAGENET dataset. After these layers, a flatten layer is used to convert the output for the dense layers. The CNN output is then passed to four distinct dense layers with $4096$, $2048$, $1024$, and $784$ neurons, all using the ReLU activation function. The second and third dense layers are preceded by a batch normalization layer. The output of the third dense layer is then passed to our Wilson-Cowan model for metapopulation, with $\mathcal{N}=784$. The total number of parameters is $129460641$, divided into $615441$ parameters for our Wilson-Cowan model for metapopulation and $128845200$ for the CNN part plus the three dense layers.
@@ A.5 IMDB
The IMDB (Internet Movie Database) dataset [73] is a comprehensive and widely-used dataset in the field of machine learning and data analysis, particularly for tasks involving natural language processing (NLP) and sentiment analysis. This dataset contains extensive information on movies, television shows, and other forms of visual entertainment. The labeled data set consists of $50000$ IMDB movie reviews, specially selected for sentiment analysis. The sentiment of reviews is binary, meaning the IMDB rating $<5$ results in a sentiment score of 00, and rating $\geq 7$ have a sentiment score of $1$. No individual movie has more than $30$ reviews. The $25000$ review labeled training set does not include any of the same movies as the $25000$ review test set. In addition, there are another $50000$ IMDB reviews provided without any rating labels.
@@
This language model is then associated with a dropout layer with a parameter $p=0.5$ and a dense layer of $512$ neurons with a sigmoid activation function. Following this, we integrate our Wilson-Cowan model for metapopulation (with $\mathcal{N}=512$), ending up with a final model with $110138626$ As stated in the main text, we perform a modification for this particular classification task. Specifically, we do not compare the final state of the dynamics with the respective planted eigenvector anymore. Instead, we apply a non-linear transformation to the final state, normalizing the inverse of equation (11), to obtain a straightforward output for our $K=2$ class problem. Consequently, we derive a probability to be in one of the two classes, allowing us to apply binary cross-entropy as the loss function and utilize binary accuracy to test the performance of our learning algorithm. We used a learning rate of $0.00003$, and we fixed the number of epochs for the fine-tuning of the whole network at $10$, with mini-batch size equal to $4$. The final time $T$ was set to be equal $4.0\Delta t^{-1}$, with $\Delta t=0.1$. For the tokenization of the dataset, we have followed the tutorial given at this link 666 https://www.tensorflow.org/text/tutorials/classify_text_with_bert .
@@ Appendix B Hardware Specification
Appendix C Visual Cortex
The visual cortex (grey, purple and green) is the primary cortical region of the brain that receives, integrates, and processes visual information relayed from the retinas. It is in the occipital lobe of the primary cerebral cortex, which is in the most posterior region of the brain. The visual cortex divides into five different areas (V1 to V5) based on function and structure. In figure 9 Panel a), only the positions of V1, V2, and V4 are presented for the sake of simplicity. Inferior Temporal (IT) is the cerebral cortex on the inferior convexity of the temporal lobe in primates including humans. It is crucial for visual object recognition and is considered to be the final stage in the ventral cortical visual system (grey and purple). The ventral stream transforms visual inputs into perceptual representations that embody the enduring characteristics of objects and their spatial relations. The ventral stream begins with V1, goes through visual area V2, then through visual area V4, and to the inferior temporal cortex. The ventral stream, is associated with form recognition, object representation and storage of long-term memory [84, 67]. The dorsal stream (grey and green) begins with V1, goes through area V2, then to the dorsomedial area and middle temporal area and to the posterior parietal cortex. The dorsal stream’s job is to mediate the visual control of skilled actions, such as reaching and grasping, directed at objects in the world. To do this, the dorsal stream needs to register visual information about the goal object on a moment-to-moment basis, transforming this information into the appropriate coordinates for the effector being used [84, 67].
Appendix D Scaling Analysis
In this appendix, we present an analysis of the scaling behavior [59] of our neural mass network model when applied to classification tasks on the MNIST and Fashion MNIST datasets. We analyze how accuracy scales with the image size. To achieve this, we create new datasets from the original ones, where the image sizes are $14\times 14$, $17\times 17$, $21\times 21$, $24\times 24$, $28\times 28$, $31\times 31$, and $35\times 35$.
Table 6 shows the accuracy as a function of the input size $\mathcal{N}$ for MNIST, while Table 7 for Fashion MNIST. For both tables, The first column identifies the size of the image, the second the value of the $\gamma$ parameter, the third one the accuracy of the model, and the last column is the time for a single epoch in seconds on our hardware B.
$\mathcal{N}$ $\gamma$ $\psi$ time for each epoch [s]
196 0.86046 0.9746 2
289 0.8822 0.98 3
441 0.8311 0.9818 4
576 1.066 0.9815 5
784 0.839 0.9813 7
961 0.827 0.9781 10
1225 1.044 0.981 12
Table 6: Scaling of neural mass network model on varying image sizes for MNIST dataset. Total number of epochs 525, batch size 200.
$\mathcal{N}$ $\gamma$ $\psi$ time for each epoch [s]
196 0.8625 0.8752 2
289 1.064 0.8794 3
441 0.971 0.8788 4
576 0.910 0.8853 5
784 0.992 0.8839 7
961 1.016 0.8847 10
1225 1.016 0.887 12
Table 7: Scaling of neural mass network model on varying image sizes for Fashion MNIST dataset. Total number of epochs 350, batch size 200.
Appendix E Ablation experiments
In this appendix, we present several ablation experiments [66] on our neural mass network model. We focus on the MNIST dataset and test the hypotheses required to enable the model to function as a classifier.
We begin by removing the stability criterion from our model while keeping all other conditions fixed. Specifically, we force the eigenvalues of the matrix $\Lambda$ to fall outside the stability region. However, when we do this, the optimization process becomes infeasible, preventing us from transforming the Wilson-Cowan model for metapopulation into a classifier.
From this reason, stability is always preserved in our model. Next, we performed ablation on the planted eigenvectors. In this setup, we maintained the stability regions for the trainable eigenvalues, retained the target structure using the fixed points of the dynamics (as presented in this work), kept the zero eigenvalues fixed, and removed the planted eigenvectors, allowing the model to learn them automatically. We observed that the model was able to construct these eigenvectors; however, it achieved a lower accuracy, classifying correctly only $97.83\%$ of the test set images after training.
We then analyzed the model under the condition that the fixed eigenvectors are not required to remain within the kernel of $\mathbf{A}$, while all other components of the model were kept unchanged. In this experiment, we set the fixed eigenvalues to $0.1$. Under these conditions, the accuracy decreased to $97.28\%$, compared to the case in which the fixed eigenvectors remain in the kernel of $\mathbf{A}$.
In conclusion, we observed that the combination of all methods used to construct the Wilson-Cowan model for metapopulation as a classifier is essential for achieving high accuracy. Furthermore, we found that the stability criterion is the most crucial component of the model; without it, our model cannot be used for classification tasks